# Quark-Gluon Plasma, Charmonia, Ultra-relativistic Heavy-Ion Collisions

### v1

@@ 4 Transverse-Momentum Spectra
In this section, we follow the methodology outlined in Refs. Zhao and Rapp (2008, 2010) to calculate charmonium $p_{\rm T}$ spectra utilizing the results from the rate equation. As discussed toward the end of Sec. 2.1, the spectra can be decomposed into primordial and regenerated components according to
@@
We solve for the primordial part by employing the Boltzmann equation without the gain term, with initial conditions obtained from pp collisions as specified in Sec. 4.1). The yield from regeneration then follows from the difference of the homogeneous solutions and the full rate equation, and we assume its $p_{\rm T}$ dependence to be given by a thermal-blastwave expression for an average regeneration temperature based on our expansion model (Sec. 4.2). This approximation has, of course, its limitations, and we will discuss evidence for that in systematic comparisons to experimental data which will be carried out in Sec. 4.3.
Figure 11: Normalized $p_{\rm T}$ spectra fitted to ALICE data of inclusive $J/\psi$ production in pp collisions at mid-rapidity Acharya et al. (2019); Aaboud et al. (2018); Sirunyan et al. (2017) (left panel), forward rapidity Acharya et al. (2023) (middle panel), and inclusive $\psi(2S)$ at forward rapidity Acharya et al. (2023).
@@ 4.1 Initial $p_{\rm T}$ spectra and their suppression in heavy-ion collisions
To construct the initial conditions of charmonia, we first perform fits to their $p_{\rm T}$ spectra in pp (5.02 TeV) collisions from the ALICE Acharya et al. (2019, 2023), ATLAS Aaboud et al. (2018) and CMS Sirunyan et al. (2017) collaborations based on the ansatz
@@
For the three parameters, we obtain $N=0.043$, $A=4.48$ and $n=3.73$ at mid-rapidity and $N=0.052$, $A=4.20$ and $n=3.90$ at forward rapidity for the $J/\psi$ spectra, and $N=0.033$, $A=5.10$ and $n=3.70$ for $\psi(2S)$ spectra at forward rapidity, see Figure 11. These spectra, subjected to nuclear-shadowing effects and with $b$-feeddown subtracted from the inclusive spectra as described in Sec. 2.4, serve as initial-momentum distributions for direct production in AA collisions at a given centrality for both forward and mid-rapidity. In addition, we account for the spatial distributions of the initial charmonia, adopting a binary-collision profile obtained from the Glauber model Miller et al. (2007), and assume a factorization between spatial and momentum distributions, $f_{\psi}(\vec{x},\vec{p},\tau_{0})=f_{\psi}(\vec{x})f_{\psi}(\vec{p})$. We are then in a position to solve the $p_{\rm T}$-dependent Boltzmann equation for the suppression of primordial production for each state, $\psi$, using the momentum-dependent rates discussed in Sec. 2.2 within our fireball model. This can be done analytically Yan et al. (2006) resulting in
@@
For inclusive spectra, as usually presented by the ALICE collaboration, we also need to add back the bottom feeddown contribution, which we assume to be conserved at the (integrated) level of 13% of the $N_{\rm coll}$-scaled yields. However, for its $p_{\rm T}$ shape in the $R_{\rm AA}$ we need to account for $b$-quark energy loss and the associated redistribution to lower momenta. This effect has been explicitly assessed in Ref. He et al. (2022) based on microscopic $b$-quark diffusion calculations He et al. (2014) and turns out to result in an approximately flat $R_{\rm AA}$ for the daughter $J/\psi$ mesons, which we will assume here for both $J/\psi$ and $\psi(2S)$.
@@ 4.2 Transverse-momentum spectra from regeneration
Regarding the $p_{\rm T}$ spectra of the regenerated component, we follow earlier works in Refs. Zhao and Rapp (2008, 2011) and employ the blastwave model based on our fireball evolution, thereby assuming that charm quarks have reached thermal equilibrium in the QGP. Based on the most recent open-charm hadron phenomenology in URHICs, which suggests $c$-quark relaxation times in the QGP of $\sim$3-4 fm He et al. (2023), and in light of the fireball lifetimes shown in Figure 6, the assumption can be justified for central AA collisions at the LHC, but is probably not quantitatively accurate in semi-central and questionable in peripheral collisions, cf. also Ref. He et al. (2022). For each charmonium state, one has
[…]
where $m_{T}=\sqrt{p_{\rm T}^{2}+m_{\psi}^{2}}$ denotes the transverse mass and $N_{0}(b)$ normalizes the absolute yield to the result of the rate equation, Eq. 1; $K_{1}$ and $I_{0}$ are the modified Bessel functions of the second and first kind, respectively. The radial flow rapidity, $\rho(r)$, is given by $\rho(r)=\tanh^{-1}\left(v_{s}\frac{r}{R}\right)$, where $R$ is the radius of the fireball and $v_{s}$ its surface velocity. We evaluate this expression at an average evolution time when most of the pertinent regeneration yield has built up, i.e., in the middle of the mixed phase for $J/\psi$ and in the hadronic phase at $T=160$ MeV for $\psi(2S)$ (e.g., $\tau$=6.6 and 9.2 fm/$c$ for central collisions, recall the left panel in Figure 7), see also Ref. Du and Rapp (2015).
Figure 12: Transverse-momentum spectra (upper panels) and pertinent nuclear-modification factors (lower panels) of inclusive $J/\psi$ production in central (left panels) and semi-central (right panels) Pb-Pb (5.02 TeV) collisions at mid-rapidity, compared to ALICE data Bai (2021); Acharya et al. (2024). The spectra in pp collisions, scaled by the pertinent binary-collision number, $N_{\rm coll}(b)$, are shown as dashed lines in the upper panels. The bands and the colors of the data have the same meaning as in Figure 8.
@@ 4.3 Comparison to experimental $p_{\rm T}$ spectra
We are now in a position to discuss our results in comparison to experimental data at the LHC, focusing on Pb-Pb(5.02 TeV) collisions. We start with the decomposition of the inclusive $J/\psi$ $p_{\rm T}$ spectra and their pertinent $R_{\rm AA}$ at mid-rapidity, as shown in Figure 12. In 0-10% central collisions (left panels), we find the well-established features of a strong suppression at high momenta, $p_{\rm T}\gtrsim 5$ GeV, and a marked rise toward lower $p_{\rm T}$ due to regeneration, reaching $R_{\rm AA}$ values of one or even larger (mostly depending on the strength of nuclear shadowing that suppresses charm production). The magnitude and shape of this bump, as well as its transition to the rather flat suppression-dominated regime at high $p_{\rm T}$, are in good agreement with ALICE data, indicating that the blastwave approximation with a collective flow of thermalized charm quarks recombining into $J/\psi$ works well. In 30-50% semi-central collisions (right panels), both the high-$p_{\rm T}$ suppression level and the low-$p_{\rm T}$ recombination bump are less pronounced; however, at the lowest $p_{\rm T}$, the data tend to be overestimated, while for intermediate $p_{\rm T}$ around $\sim$5 GeV, the data are underestimated. This discrepancy indicates that the assumption of a thermalized blastwave for the recombining charm quarks is not accurate anymore; indeed, Refs. He et al. (2022); Du and Rapp (2022) have shown that employing transported $c$-quark spectra, which do not fully thermalize in semi-central collisions, remedies this discrepancy. In particular, the crossing between the regeneration and the primordial contribution will be shifted to higher $p_{\rm T}$, closer to 5 GeV in semi-central collisions, rather than 3 GeV as implied by the thermalized blastwave approximation depicted in the right panels of Figure 12.
@@
Figure 14: Forward-rapidity $J/\psi$ $p_{\rm T}$ spectra (upper panels) and $R_{\rm AA}$ (lower panels) compared to ALICE data Acharya et al. (2020a, 2024) for 3 centrality selections. Bands and lines have the same meaning as in Figure 8.
Our predictions for the $p_{\rm T}$ dependence of $\psi(2S)$ production at mid-rapidity are shown in Figure 13. At the level of the $R_{\rm AA}$, the regeneration is less prominent for $\psi(2S)$ than for $J/\psi$. The later production in the time evolution of the fireball leads to a significant shift of the maximum of the “flow bump" out to higher $p_{\rm T}$ compared to $J/\psi$, as a direct consequence of the “sequential regeneration" Du and Rapp (2015). This effect is also visible when comparing central to semi-central collisions.
@@
Figure 16: Nuclear-modification factor as a function of $p_{\rm T}$ for inclusive $J/\psi$ (blue) and $\psi(2S)$ (red) production at forward rapidity in 0-90% Pb-Pb(5.02 TeV) collisions, compared to ALICE data Acharya et al. (2024).
@@
Experimental data for the $p_{\rm T}$ dependence of $\psi(2S)$ production in 0-90% (essentially “minimum bias") Pb-Pb collisions have recently become available in Ref. Acharya et al. (2024), which also contains our original predictions (consistent with our calculations presented herein). The pertinent $R_{\rm AA}(p_{\rm T})$ is illustrated in Figure 16, alongside the $J/\psi$ results. Note that the 0-90% centrality selection exhibits a rather strong bias toward central collisions, due to the approximate $N_{\rm coll}$ scaling of the hard-produced $c\bar{c}$ pairs (both open and hidden). Consequently, there is only a slight indication that the blastwave approximation for regenerated $J/\psi$ is inaccurate, as it under-predicts the data around $p_{\rm T}\simeq$ 5 GeV. For the $\psi(2S)$, the predictions work out well, attesting to the predictive power of our transport framework. Recall that $\psi(2S)$ recombination occurs later in the fireball’s evolution than that of $J/\psi$, allowing more time for the open-charm particles to relax toward equilibrium; one furthermore finds the “flow bump" to be moved out to higher $p_{\rm T}$ compared to the $J/\psi$, which is not conclusive from the data (yet).
Refer to captionRefer to caption
@@
Figure 17: Transverse-momentum dependent $\psi(2S)$ over $J/\psi$ ratio (left panel) and their $R_{\rm AA}$ double ratio (right panel) at forward rapidity compared to ALICE data Acharya et al. (2024).
@@
Figure 18: Inclusive $J/\psi$ nuclear-modification factor as a function of $N_{\rm part}$ for Pb-Pb collisions at 5.02 TeV for four $p_{\rm T}$ intervals, compared with data Acharya et al. (2020a)
@@
The $p_{\rm T}$ dependent ratio of $\psi(2S)$ to $J/\psi$ yields at forward rapidity, as depicted in the left panel of Figure 17, also aligns reasonably well with the ALICE data Acharya et al. (2024). The increasing trend with $p_{\rm T}$ from our fits to pp data tends to overestimate the data for the ratio at the highest $p_{\rm T}$. This discrepancy also migrates into the AA result (which may be partly due to the underestimation of the $J/\psi$ production at intermediate $p_{\rm T}$). Indeed, the agreement is better when dividing out the pp reference spectra in the $R_{\rm AA}$ double ratio, shown in the right panel of Figure 17. Here, the slight deficit in the $J/\psi$ yield at intermediate $p_{\rm T}$ shows up as a mild maximum structure which is not observed in the data at this point. In principle, such a maximum could be another signature of sequential regeneration in central collisions, when the open-charm spectra are close to thermal equilibrium.
@@
We further investigate the $R_{\rm AA}(N_{\rm part})$ for inclusive $J/\psi$ production at forward rapidity, presented in Figure 14 but binned into different $p_{\rm T}$ intervals. The corresponding results, alongside ALICE data Acharya et al. (2020a), are depicted in Figure 18. The $p_{\rm T}$-dependent shadowing effect, concentrated at low $p_{\rm T}$, introduces an uncertainty in the $R_{\rm AA}$ which is most pronounced at low $p_{\rm T}$ but diminishes at higher $p_{\rm T}$. In peripheral collisions, the nuclear modification factor is primarily influenced by the primordial contributions across all $p_{\rm T}$ regions, approaching 1 for small $N_{\rm part}$. Conversely, central collisions the scenario changes. For the lower $p_{\rm T}$ bins, regeneration processes predominantly contribute to the observed $J/\psi$ production, leading to an increasing $R_{\rm AA}$, while at higher $p_{\rm T}$, the contributions shift towards non-prompt and primordial $J/\psi$ production mechanisms, rendering a characteristic decrease in $R_{\rm AA}$ values with increasing $N_{\rm part}$.
@@
Figure 19: $\langle p_{\rm T}\rangle$ (upper left) and $\langle p_{\rm T}^{2}\rangle$ (upper right) of $J/\psi$ in 5.02 TeV pp and Pb-Pb collisions at mid-rapidity. Their ratios relative to the pp results are shown in the lower panels; ALICE data are from Refs. Acharya et al. (2024, 2020b).
@@
Finally, we attend to the centrality dependence of the average charmonium momentum, $\langle p_{\rm T}\rangle$, and its square, $\langle p_{\rm T}^{2}\rangle$. We calculate a “small" nuclear-modification factor" defined in Ref. Zhou et al. (2014), normalized to the values in pp and Pb-Pb collisions,
@@
the results are shown in Figure 19. In central collisions, the predominance of regenerated $J/\psi$’s, which typically exhibit a softer spectrum, leads to a reduction in both the transverse momentum and its square, causing their $r_{\rm AA}$’s to fall below 1. This effect is more pronounced for the latter and also shows a more gradual fall-off with centrality. Our model calculations describe the centrality dependence of $\langle p_{\rm T}\rangle$ rather well, yet slightly underestimate the $\langle p_{\rm T}^{2}\rangle$ data in semi-central collisions. Again, this can be traced back to a recombination contribution that is too soft in semi-central collisions, most likely since charm quarks do not achieve full thermal equilibrium, which can be remedied by employing explicitly transported charm-quark distributions He et al. (2022); Du and Rapp (2022).

### v2

@@ 4 Transverse-Momentum Spectra
In this section, we follow the methodology outlined in Refs. Zhao:2010nk to calculate charmonium $p_{\rm T}$ spectra utilizing the results from the rate equation. As discussed toward the end of Sec. 2.1, the spectra can be decomposed into primordial and regenerated components according to
@@
We solve for the primordial part by employing the Boltzmann equation without the gain term, with initial conditions obtained from $p$ collisions as specified in Sec. 4.1). The yield from regeneration then follows from the difference of the homogeneous solutions and the full rate equation, and we assume its $p_{\rm T}$ dependence to be given by a thermal-blastwave expression for an average regeneration temperature based on our expansion model (Sec. 4.2). This approximation has, of course, its limitations, and we will discuss evidence for that in systematic comparisons to experimental data which will be carried out in Sec. 4.3.
@@ 4.1 Initial $p_{\rm T}$ spectra and their suppression in heavy-ion collisions
Figure 11: Normalized $p_{\rm T}$ spectra fitted to ALICE data of inclusive $J/\psi$ production in $p$ collisions at mid-rapidity ALICE:2019pid; ATLAS:2017prf; CMS:2017exb (left panel), forward rapidity ALICE:2021qlw (middle panel), and inclusive $\psi(2S)$ at forward rapidity ALICE:2021qlw.
To construct the initial conditions of charmonia, we first perform fits to their $p_{\rm T}$ spectra in $p$ (5.02 TeV) collisions from the ALICE ALICE:2019pid; ALICE:2021qlw, ATLAS ATLAS:2017prf and CMS CMS:2017exb collaborations based on the ansatz
@@
For the three parameters, we obtain $N=0.043$, $A=4.48$ and $n=3.73$ at mid-rapidity and $N=0.052$, $A=4.20$ and $n=3.90$ at forward rapidity for the $J/\psi$ spectra, and $N=0.033$, $A=5.10$ and $n=3.70$ for $\psi(2S)$ spectra at forward rapidity, see Figure 11. These spectra, subjected to nuclear-shadowing effects and with $b$-feeddown subtracted from the inclusive spectra as described in Sec. 2.4, serve as initial-momentum distributions for direct production in AA collisions at a given centrality for both forward and mid-rapidity. In addition, we account for the spatial distributions of the initial charmonia, adopting a binary-collision profile obtained from the Glauber model Miller:2007ri, and assume a factorization between spatial and momentum distributions, $f_{\psi}(\vec{x},\vec{p},\tau_{0})=f_{\psi}(\vec{x})f_{\psi}(\vec{p})$. We are then in a position to solve the $p_{\rm T}$-dependent Boltzmann equation for the suppression of primordial production for each state, $\psi$, using the momentum-dependent rates discussed in Sec. 2.2 within our fireball model. This can be done analytically Yan:2006ve resulting in
@@
For inclusive spectra, as usually presented by the ALICE collaboration, we also need to add back the bottom feeddown contribution, which we assume to be conserved at the (integrated) level of 13% of the $N_{\rm coll}$-scaled yields. However, for its $p_{\rm T}$ shape in the $R_{\rm AA}$ we need to account for $b$-quark energy loss and the associated redistribution to lower momenta. This effect has been explicitly assessed in Ref. He:2021zej based on microscopic $b$-quark diffusion calculations and turns out to result in an approximately flat $R_{\rm AA}$ for the daughter $J/\psi$ mesons, which we will assume here for both $J/\psi$ and $\psi(2S)$.
@@ 4.2 Transverse-momentum spectra from regeneration
Regarding the $p_{\rm T}$ spectra of the regenerated component, we follow earlier works in Refs. Zhao:2011cv and employ the blastwave model based on our fireball evolution, thereby assuming that charm quarks have reached thermal equilibrium in the QGP. Based on the most recent open-charm hadron phenomenology in URHICs, which suggests $c$-quark relaxation times in the QGP of $\sim$3-4 fm He:2022ywp; Capellino:2022nvf, and in light of the fireball lifetimes shown in Figure 6, the assumption can be justified for central AA collisions at the LHC, but is probably not quantitatively accurate in semi-central and questionable in peripheral collisions, cf. also Ref. He:2021zej. For each charmonium state, one has
$\frac{\mathrm{d}N^{\rm reg}_{\psi}}{\mathrm{d}p_{T}^{2}}=N_{0}(b)m_{T}\int_{0}^{R}rdrK_{1}(\frac{m_{T}\cosh\rho(r)}{T})I_{0}(\frac{p_{\rm T}\sinh\rho(r)}{T})\ ,$ (19)
where $m_{T}=\sqrt{p_{\rm T}^{2}+m_{\psi}^{2}}$ denotes the transverse mass and $N_{0}(b)$ normalizes the absolute yield to the result of the rate equation, Eq. 1; $K_{1}$ and $I_{0}$ are the modified Bessel functions of the second and first kind, respectively. The radial flow rapidity, $\rho(r)$, is given by $\rho(r)=\tanh^{-1}\left(v_{s}\frac{r}{R}\right)$, where $R$ is the radius of the fireball and $v_{s}$ its surface velocity. We evaluate this expression at an average evolution time when most of the pertinent regeneration yield has built up, i.e., in the middle of the mixed phase for $J/\psi$ and in the hadronic phase at $T=160$ MeV for $\psi(2S)$, independent of centrality (e.g., $\tau$=6.6 and 9.2 fm/$c$ for central collisions, recall the left panel in Figure 7), see also Ref. Du:2015wha.
Figure 12: Transverse-momentum spectra (upper panels) and pertinent nuclear-modification factors (lower panels) of inclusive $J/\psi$ production in central (left panels) and semi-central (right panels) Pb-Pb (5.02 TeV) collisions at mid-rapidity, compared to ALICE data Bai:2020svs; ALICE:2023gco. The spectra in $p$ collisions, scaled by the pertinent binary-collision number, $N_{\rm coll}(b)$, are shown as dashed lines in the upper panels. The bands and the colors of the data have the same meaning as in Figure 8.
@@ 4.3 Comparison to experimental $p_{\rm T}$ spectra
We are now in a position to discuss our results in comparison to experimental data at the LHC, focusing on Pb-Pb(5.02 TeV) collisions. We start with the decomposition of the inclusive $J/\psi$ $p_{\rm T}$ spectra and their pertinent $R_{\rm AA}$ at mid-rapidity, as shown in Figure 12. In 0-10% central collisions (left panels), we find the well-established features of a strong suppression at high momenta, $p_{\rm T}\gtrsim 5\,{\rm GeV}/c$, and a marked rise toward lower $p_{\rm T}$ due to regeneration, reaching $R_{\rm AA}$ values of one or even larger (mostly depending on the strength of nuclear shadowing that suppresses charm production). The magnitude and shape of this bump, as well as its transition to the rather flat suppression-dominated regime at high $p_{\rm T}$, are in good agreement with ALICE data, indicating that the blastwave approximation with a collective flow of thermalized charm quarks recombining into $J/\psi$ works well. In 30-50% semi-central collisions (right panels), both the high-$p_{\rm T}$ suppression level and the low-$p_{\rm T}$ recombination bump are less pronounced; however, at the lowest $p_{\rm T}$, the data tend to be overestimated, while for intermediate $p_{\rm T}$ around $\sim 5\,{\rm GeV}/c$, the data are underestimated. This discrepancy indicates that the assumption of a thermalized blastwave for the recombining charm quarks is not accurate anymore; indeed, Refs. He:2021zej; Du:2022uvj have shown that employing transported $c$-quark spectra, which do not fully thermalize in semi-central collisions, remedies this discrepancy. In particular, the crossing between the regeneration and the primordial contribution will be shifted to higher $p_{\rm T}$, closer to $5\,{\rm GeV}/c$ in semi-central collisions, rather than $3\,{\rm GeV}/c$ as implied by the thermalized blastwave approximation depicted in the right panels of Figure 12.
Figure 14: Forward-rapidity $J/\psi$ $p_{\rm T}$ spectra (upper panels) and $R_{\rm AA}$ (lower panels) compared to ALICE data ALICE:2019lga; ALICE:2023gco for 3 centrality selections. Bands and lines have the same meaning as in Figure 8.
Our predictions for the $p_{\rm T}$ dependence of $\psi(2S)$ production at mid-rapidity are shown in Figure 13. At the level of the $R_{\rm AA}$, the regeneration is less prominent for $\psi(2S)$ than for $J/\psi$. The later production in the time evolution of the fireball leads to a significant shift of the maximum of the “flow bump" out to higher $p_{\rm T}$ compared to $J/\psi$, as a direct consequence of the “sequential regeneration" Du:2015wha. This effect is also visible when comparing central to semi-central collisions.
@@
Figure 16: Nuclear-modification factor as a function of $p_{\rm T}$ for inclusive $J/\psi$ (blue) and $\psi(2S)$ (red) production at forward rapidity in 0-90% Pb-Pb(5.02 TeV) collisions, compared to ALICE data ALICE:2022jeh.
@@
Figure 17: Transverse-momentum dependent $\psi(2S)$ over $J/\psi$ ratio (left panel) and their $R_{\rm AA}$ double ratio (right panel) at forward rapidity compared to ALICE data ALICE:2022jeh.
@@
Figure 18: Inclusive $J/\psi$ nuclear-modification factor as a function of $N_{\rm part}$ for Pb-Pb collisions at 5.02 TeV for four $p_{\rm T}$ intervals, compared with data ALICE:2019lga
Experimental data for the $p_{\rm T}$ dependence of $\psi(2S)$ production in 0-90% (essentially “minimum bias") Pb-Pb collisions have recently become available in Ref. ALICE:2022jeh, which also contains our original predictions (consistent with our calculations presented herein). The pertinent $R_{\rm AA}(p_{\rm T})$ is illustrated in Figure 16, alongside the $J/\psi$ results. Note that the 0-90% centrality selection exhibits a rather strong bias toward central collisions, due to the approximate $N_{\rm coll}$ scaling of the hard-produced $c\bar{c}$ pairs (both open and hidden). Consequently, there is only a slight indication that the blastwave approximation for regenerated $J/\psi$ is inaccurate, as it under-predicts the data around $p_{\rm T}\simeq 5\,{\rm GeV}/c$. For the $\psi(2S)$, the predictions work out well, attesting to the predictive power of our transport framework. Recall that $\psi(2S)$ recombination occurs later in the fireball’s evolution than that of $J/\psi$, allowing more time for the open-charm particles to relax toward equilibrium; one furthermore finds the “flow bump" to be moved out to higher $p_{\rm T}$ compared to the $J/\psi$, which is not conclusive from the data (yet).
@@
The $p_{\rm T}$ dependent ratio of $\psi(2S)$ to $J/\psi$ yields at forward rapidity, as depicted in the left panel of Figure 17, also aligns reasonably well with the ALICE data ALICE:2022jeh. The increasing trend with $p_{\rm T}$ from our fits to $p$ data tends to overestimate the data for the ratio at the highest $p_{\rm T}$. This discrepancy also migrates into the AA result (which may be partly due to the underestimation of the $J/\psi$ production at intermediate $p_{\rm T}$). Indeed, the agreement is better when dividing out the $p$ reference spectra in the $R_{\rm AA}$ double ratio, shown in the right panel of Figure 17. Here, the slight deficit in the $J/\psi$ yield at intermediate $p_{\rm T}$ shows up as a mild maximum structure which is not observed in the data at this point. In principle, such a maximum could be another signature of sequential regeneration in central collisions, when the open-charm spectra are close to thermal equilibrium.
@@
We further investigate the $R_{\rm AA}(N_{\rm part})$ for inclusive $J/\psi$ production at forward rapidity, presented in Figure 14 but binned into different $p_{\rm T}$ intervals. The corresponding results, alongside ALICE data ALICE:2019lga, are depicted in Figure 18. The $p_{\rm T}$-dependent shadowing effect, concentrated at low $p_{\rm T}$, introduces an uncertainty in the $R_{\rm AA}$ which is most pronounced at low $p_{\rm T}$ but diminishes at higher $p_{\rm T}$. In peripheral collisions, the nuclear modification factor is primarily influenced by the primordial contributions across all $p_{\rm T}$ regions, approaching 1 for small $N_{\rm part}$. Conversely, central collisions the scenario changes. For the lower $p_{\rm T}$ bins, regeneration processes predominantly contribute to the observed $J/\psi$ production, leading to an increasing $R_{\rm AA}$, while at higher $p_{\rm T}$, the contributions shift towards non-prompt and primordial $J/\psi$ production mechanisms, rendering a characteristic decrease in $R_{\rm AA}$ values with increasing $N_{\rm part}$.
@@
Figure 19: $\langle p_{\rm T}\rangle$ (upper left) and $\langle p_{\rm T}^{2}\rangle$ (upper right) of $J/\psi$ in 5.02 TeV $p$ and Pb-Pb collisions at mid-rapidity. Their ratios relative to the $p$ results are shown in the lower panels; ALICE data are from Refs. ALICE:2022jeh; ALICE:2019nrq.
@@
Finally, we attend to the centrality dependence of the average charmonium momentum, $\langle p_{\rm T}\rangle$, and its square, $\langle p_{\rm T}^{2}\rangle$. We calculate a “small" nuclear-modification factor" defined in Ref. Zhou:2014kka, normalized to the values in $p$ and Pb-Pb collisions,
@@
the results are shown in Figure 19. In central collisions, the predominance of regenerated $J/\psi$’s, which typically exhibit a softer spectrum, leads to a reduction in both the transverse momentum and its square, causing their $r_{\rm AA}$’s to fall below 1. This effect is more pronounced for the latter and also shows a more gradual fall-off with centrality. Our model calculations describe the centrality dependence of $\langle p_{\rm T}\rangle$ rather well, yet slightly underestimate the $\langle p_{\rm T}^{2}\rangle$ data in semi-central collisions. Again, this can be traced back to a recombination contribution that is too soft in semi-central collisions, most likely since charm quarks do not achieve full thermal equilibrium, which can be remedied by employing explicitly transported charm-quark distributions He:2021zej; Du:2022uvj.